OpenStudy (anonymous):

MEDAL + FAN! :D 3y^2-6y+x+5=0 How to set it to vertex form?

OpenStudy (anonymous):

pls :'(

OpenStudy (anonymous):

pllllllllllllllllllllllllllllllllllllllllllssssssssssssssssssssssssssssssssssss

OpenStudy (cwrw238):

first take 3 out as follows 3(y^2 - 2y) + x +5 = 0 3[(y - 1)^2 - 1)] + x + 5 = 0 3(y -1)^2 - 3 + x + 5 = 0 3(y - 1)^2 + x + 2 = 0

OpenStudy (anonymous):

Sorry for the late response and thanks.